# Understanding and Mitigating Membership Inference Risks of Neural Ordinary Differential Equations

### v1

@@ 3.4. Characterizing the Membership Risks
We now run an in-depth analysis of the factors that may attribute to the membership risks. We first examine the choice of ODE solvers (a block-level configurations) and then analyze model-level factors: the number of ODE blocks and the choice of equations used for modeling a system. Figure 2 summarizes our results.
@@
Figure 2. Characterization of membership risks in NODEs with different configurations. We vary the choice of ODE solvers (left), the number of ODE blocks (middle), and the equations used for modeling a system (right). We show the results in CIFAR-10 against LiRA (Carlini et al., 2022), the strongest membership inference attack. Please refer to Appendix for details and results in CIFAR-100.
@@
Impact of a solver. In all our experiments, we use Euler’s method as the default solver; but, other popular choices exist in the literature. We therefore analyze how the choice of solver impacts the membership risks of NODEs. We train and test the same NODEs with two other solvers: a fixed-step solver, Runge–Kutta (RK4) (Butcher, 2016) and an adaptive-step solver, Dormand–Prince (Dopri5) (Dormand and Prince, 1980).
@@
The leftmost figure illustrates the success rate of LiRA on NODEs trained and tested with three different solvers. Overall, we observe a marginal difference in attack success across these models. In both CIFAR-10 and CIFAR-100 (shown in Appendix C), the TPR remains 0.7–1.0% and 3.0–3.3% @ 0.1% FPR, respectively. Upon closer examination, Euler and RK4 yield similar risks, while Dopri5 shows $\sim$0.3lower TPR at the same FPR. We attribute this to Dopri5’s use of adaptive step-sizes. During the training of a NODE model with Dopri5, it stops processing of an input when the error rate becomes sufficiently small–within the predefined tolerance value. It acts as an early-stopping mechanism (Song and Mittal, 2021), preventing the model from becoming overconfident on specific training samples. On the other hand, RK4, by modeling data with higher-order polynomials may achieve an increased training (and testing) accuracy. However, the ability to perform complex modeling can introduce a risk of overfitting to specific training samples, which can also increase membership inference risks. In CIFAR-100, a more complex task than CIFAR-10, we observe 0.3% and 1.0% higher TPR at 0.1% and 1% FPR, respectively, compared to the model using Euler.
@@
Impact of the number of ODE blocks. Our NODEs replace a group of residual blocks with an ODE block, a configuration chosen for its comparable accuracy to ResNet models. But we can increase the number of blocks further. We hypothesize that, because additional ODE blocks enhance the model’s ability to represent complex systems, this may lead to an increase in membership risks.
@@
The middle figure shows the membership risks of NODEs when the number of ODE blocks is increased from 1 to 3. We observe that increasing the number of ODE blocks leads to higher membership risks. This is primarily because more parameters increase the likelihood of overfitting. However, we do not observe significant performance gains from the decision, meaning that simply adding blocks is not an ideal strategy for improving the performance.
@@
Non-stochastic NODE variants. In Sec. 2.1, we review NODE variants designed to enhance expressivity and stability, ultimately improving generalization. We examine whether these benefits come with a reduction in membership risks. We evaluate ANODE (Dupont et al., 2019a), SONODE (Norcliffe et al., 2020) and HBNODE (Xia et al., 2021). ANODE introduces additional dimensions to NODEs, and SONODE extends the modeling equations to second order, while HBNODE incorporates momentum into the system equations, stabilizing the ODEs learned by the model. We select ODENet-64 as our baseline model because it is composed of a single ODE block, making it easy to extend to NODE variants. We implement SONODE and HBNODE using an ODE block identical to that in ODENet-64. We augment ODENet-64 with an additional 16 and 64 dimensions (ANODE). These variants are expected to enhance the model’s generalization capabilities, allowing us to assess the impact of better system modeling on membership risks.
@@
The rightmost figure compares the membership risks of these models with those of the baseline ODENet-64 model. We first observe that NODE variants, except for the HBNODE model, achieves higher test accuracy (84.6%–86.5%) than the baseline (84.4%). SONODE performs the best with 86.5%, followed by ANODE (+64) at 85.1% and ANODE (+16) at 84.6%. But HBNODE achieves 82.1%. Interestingly, we find that all the NODE variants are less vulnerable to LiRA. SONODE, ANODE (+64), ANODE (+16), and HBNODE achieve TPRs of 0.44%, 0.35%, 0.24%, and 0.24% at 0.1% FPR, respectively, compared to ODENet-64 with a TPR of 0.86% at 0.1% FPR. We attribute this to the reduced overfitting shown by NODE variants. The differences between training and testing accuracy are 5.9%, 4.0%, 3.0%, and 2.8% for SONODE, ANODE (+64), ANODE (+16), and HBNODE, respectively, compared to 9.7% gap in ODENet-64. However, for HBNODE, there is a possibility that the lower risk is attributed to the lack of generalization. Modeling NODEs with momentum (HBNODE) enhances the stability, but does not increase a model’s capacity. In our evaluation, the training accuracy of HBNODE is 84.9%, 2.8–7% lower than that of the other variants.
4. NSDEs Are DP-Learners
@@
In the previous section, we demonstrated that NODEs and their non-stochastic variants are still vulnerable to membership inference attacks. We now shift our focus to the stochastic variants of NODEs, particularly neural stochastic differential equations (NSDEs) (Liu et al., 2020). In NSDEs, which contain “drift” and “diffusion” terms, stochasticity arises from the randomness in the diffusion term, typically modeled as white or colored Gaussian noise. We hypothesize that this diffusion term leads NSDEs to function as a differentially-private (DP) mechanism (Dwork, 2006), preserving the privacy of its inputs along its evolution. Here we first formally show that NSDEs are DP-learners. Because of the diffusion term, those models learn “differentially-private” systems of ODEs, achieving privacy without the need of additional mechanisms, such as DP-SGD (Abadi et al., 2016), which alters the training process. Next in Sec 5, we empirically assess the membership risks of NSDEs in comparison to non-stochastic NODE variants and “private” feedforward models trained with DP-SGD.
4.1. Theoretical Analysis
@@
in terms of a drift function ${}f:\Omega\to\mathbb{R}^{n}$ and a diffusion term ${}G:\Omega\to\mathbb{R}^{n\times n}$, where the explicit dependence on parameters $\Theta$ has been dropped for notational convenience. Our privacy bounds will make use of the sensitivity of a function ${}g:\Omega\to\mathbb{R}^{n}$ (c.f. (Dwork, 2006)):
@@
Here, we adopt the following natural assumptions from the work by Liu et al. (Liu et al., 2020), which guarantees the well-posedness of the SDE formulation under consideration.
@@
Theorem 4.1 (SDEs Are Differentially-Private).
@@
For any privacy budget $10<\varepsilon<$ 1 and failure probability $\delta>0$, the mechanism ${}M(~{}h)=~{}h(T)$, defined by the flow map of the SDE at time $T$, is $(\varepsilon,\delta)$-differentially private, provided that $\sigma\geq\sqrt{2\ln(1.25/\delta)}(TL_{~{}f}/\varepsilon)$ [^2] Note that $\varepsilon$ and $\delta$ here are different from those used in Sec 2.3..
@@
where ${}B_{T}\sim\mathcal{N}\left(~{}0,\sigma^{2}~{}I\right)$ represents additive Gaussian noise and ${}F$ denotes deterministic part of the SDE evolution. Proposition B.1 in Appendix B shows that this mechanism obeys $(\varepsilon,\delta)$-differential privacy whenever $\varepsilon<1$ and $\sigma\geq\sqrt{2\ln(1.25/\delta)}(S_{~{}F}/\varepsilon)$. The claimed bound on the variance $\sigma$ follows since the sensitivity $S_{~{}F}$ is bounded above due to Lipschitz continuity in ${}f$:
@@
Corollary 4.2 (NSDE Training is Differentially Private).
Consider target privacy parameters $10\!<\!\varepsilon^{\prime}\!<\!$ 1 and $\delta^{\prime}\!>\!0$ with a maximum iteration count of $K$. Then, NSDE training of the drift ${}f$ in Theorem 4.1 will be $(\varepsilon^{\prime},K\delta+\delta^{\prime})$-differentially private as long as the variance satisfies $\sigma\!\geq\!4\sqrt{K\ln(1.25/\delta)\ln(1/\delta^{\prime})}(TL/\varepsilon^{\prime})$. Here, $\delta\!>\!0$ is arbitrary and $L=\max_{~{}f}L_{~{}f}$ denotes the maximum Lipschitz constant attainable by ${}f$ during training.
@@
Note that, by the parallel composition principle, mini-batching into disjoint subsets at each iteration does not increase the total privacy leakage, as each individual’s data in the state ${}h$ is accessed only once by the SDE mechanism. Similarly, the composition of the NSDE output with a downstream classifier does not degrade $(\varepsilon,\delta)$-differential privacy guarantees, as these guarantees are immune to post-processing by construction. Consequently, the total privacy leakage in NSDE training is directly proportional to the number of training iterations $K$, corresponding to the number of accesses to the private input ${}h$. The result follows from the Strong Composition Theorem in Dwork et al. (Dwork and Roth, 2014) [Theorem 3.20, Corollary 3.21], in combination with the variance bound established in Theorem 4.1. ∎
@@
Note that determining the Lipschitz constant $L$ in Corollary 4.2 is most easily accomplished by restricting the NSDE architecture to guarantee a particular Lipschitz bound, see, e.g., (Wang and Manchester, 2023). Moreover, it should be noted that the privacy guarantees presented here are not necessarily tight in every case and may be improved with more sophisticated privacy accounting mechanisms, such as those in (Abadi et al., 2016; Dong et al., 2022).
@@
Corollary 4.2 guarantees the privacy of NSDE training provided that the diffusion term is constructed to add enough stochastic noise along NSDE solution trajectories. This has the remarkable advantage of privatizing learning without requiring direct intervention in the training process: the network dynamics are already private, therefore no further privacy-preserving modifications are required during the backward pass. In contrast, DP-SGD must use inexact gradient information in order to retain privacy, which can potentially hinder the training process.
@@
Table 2. Contrasting membership inference risks of NSDEs and NODEs. We evaluate both models with five existing attacks.
@@
Acc. (C-10) Acc. (C-100) TPR @ 0.1% FPR TPR @ 1% FPR AUC Inference acc. Model Method Train Test Train Test C-10 C-100 C-10 C-100 C-10 C-100 C-10 C-100 NODE (ODENet-16-32-64) Yeom et al. (Yeom et al., 2018) 94.3% 84.4% 68.2% 52.4% 0.00% 0.08% 0.00% 1.11% 0.558 0.626 56.59% 61.01% Shokri et al. (Shokri et al., 2017) 0.04% 0.06% 0.43% 0.66% 0.448 0.449 50.01% 50.01% Song and Mittal (Song et al., 2021) 0.00% 0.13% 0.00% 1.43% 0.529 0.498 54.22% 51.37% Watson et al. (Watson et al., 2022) 0.00% 0.08% 0.00% 1.11% 0.558 0.626 56.69% 61.01% Carlini et al. (Carlini et al., 2022) 1.01% 1.10% 4.22% 5.33% 0.616 0.676 57.65% 62.52% NSDE (SDENet $k\!=\!2$) Yeom et al. (Yeom et al., 2018) 81.9% 81.9% 56.5% 50.6% 0.02% 0.05% 0.92% 1.09% 0.525 0.588 52.44% 56.96% Shokri et al. (Shokri et al., 2017) 0.09% 0.08% 0.86% 0.85% 0.484 0.473 50.04% 50.00% Song and Mittal (Song et al., 2021) 0.00% 0.12% 0.20% 1.34% 0.509 0.498 50.95% 50.84% Watson et al. (Watson et al., 2022) 0.02% 0.05% 0.91% 1.09% 0.525 0.588 52.44% 56.79% Carlini et al. (Carlini et al., 2022) 0.14% 0.43% 1.24% 3.00% 0.525 0.612 51.89% 57.64%
5. Empirical Evaluation

### v2

@@ 3.4 Characterizing the Membership Risks
We now analyze factors that attribute to the membership inference risks. For this analysis, we use LiRA, where NODEs demonstrate the largest reduction in their risk. We examine block-level configurations, such as the choice of ODE solvers, and model-level factors, including the number of ODE blocks and the type of differential equations used to model a system. Figure 2 summarizes our results.
@@
Figure 2: Characterization of membership risks in NODEs with different configurations. We vary the choice of ODE solvers (left), the number of ODE blocks (middle), and the equations used for modeling a system (right). We show the results in CIFAR-10 against LiRA [6], the strongest membership inference attack. Please refer to Appendix for details and results in CIFAR-100.
@@
Impact of a solver. In all our experiments, we use Euler’s method as the default solver; but, other popular choices exist in the literature. We therefore analyze how the choice of solver impacts the membership risks of NODEs. We train and test the same NODEs with two other solvers: a fixed-step solver, Runge–Kutta (RK4) and an adaptive-step solver, Dormand–Prince (Dopri5).
@@
The leftmost figure illustrates the success rate of LiRA on NODEs trained and tested with three different solvers. Overall, we observe a marginal difference in attack success across these models. In both CIFAR-10 and CIFAR-100 (shown in Appendix), the TPR remains 0.7–1.0% and 3.0–3.7% @ 0.1% FPR, respectively. Upon closer examination, Euler and RK4 yield similar risks, while Dopri5 shows 0.3% lower TPR at the same FPR. We attribute this to Dopri5’s use of adaptive step-sizes. During the training of a NODE model with Dopri5, it stops processing of an input when the error rate becomes sufficiently small–within the predefined tolerance value. It acts as an early-stopping mechanism [5], preventing the model from becoming overconfident on specific training samples. RK4, by modeling data with higher-order polynomials, may achieve an increased training (and testing) accuracy. However, the ability to perform complex modeling can introduce a risk of overfitting to specific training samples, which can also increase membership inference risks. In CIFAR-100, a more complex task than CIFAR-10, we observe 0.3% and 1.0% higher TPR at 0.1% and 1% FPR, respectively, compared to the model using Euler.
@@
Impact of the number of ODE blocks. Our NODEs replace a group of residual blocks with an ODE block, a configuration chosen for its comparable accuracy to ResNet models. One can also increase the number of ODE blocks further. We hypothesize that, because additional ODE blocks enhance the model’s ability to represent complex systems, this may lead to an increase in membership risks.
@@
The middle figure compares the membership risks of NODEs as the number of ODE blocks increases from 1 to 3. We find that adding more ODE blocks leads to higher membership risks, likely due to the increased number of parameters, which raises the likelihood of overfitting. However, this increase does not yield significant performance gains, suggesting that simply adding more ODE blocks is not an effective strategy for improving the performance.
@@
Non-stochastic NODE variants. In §2.1, we review NODE variants designed to enhance expressivity and stability, ultimately improving generalization. We examine whether these benefits come with a reduction in membership risks. We evaluate ANODE [16], SONODE [21] and HBNODE [18]. We select ODENet-64 as our baseline model because it is composed of a single ODE block, making it easy to extend to NODE variants. We implement SONODE and HBNODE using an ODE block identical to that in ODENet-64. We augment ODENet-64 with an additional 16 and 64 dimensions (ANODE). These variants are expected to enhance the model’s generalization capabilities, allowing us to assess the impact of better system modeling on membership risks.
@@
The rightmost figure compares the membership risks of these models with those of the baseline ODENet-64 model. We first observe that NODE variants, except for the HBNODE model, achieves higher test accuracy (84.6–86.5%) than the baseline (84.4%). SONODE performs the best with 86.5%, followed by ANODE (+64) at 85.1% and ANODE (+16) at 84.6%. But HBNODE achieves 82.1%. Interestingly, we find that all the NODE variants are less vulnerable to LiRA. SONODE, ANODE (+64), ANODE (+16), and HBNODE achieve TPRs of 0.44%, 0.35%, 0.24%, and 0.24% at 0.1% FPR, respectively, compared to ODENet-64 with a TPR of 0.86% at 0.1% FPR. We attribute this to the reduced overfitting shown by NODE variants. The differences between training and testing accuracy are 5.9%, 4.0%, 3.0%, and 2.8% for SONODE, ANODE (+64), ANODE (+16), and HBNODE, respectively, compared to 9.7% gap in ODENet-64. However, for HBNODE, there is a possibility that the lower risk is attributed to the lack of generalization. Modeling NODEs with momentum (HBNODE) enhances the stability, but does not increase a model’s capacity. In our evaluation, the training accuracy of HBNODE is 84.9%, 2.8–7% lower than that of the other variants.
4 NSDEs Are DP-Learners
@@
In the previous section, we demonstrated that NODEs and their non-stochastic variants are still vulnerable to membership inference attacks. We now shift our focus to the stochastic variants of NODEs, particularly neural stochastic differential equations (NSDEs) [32]. In NSDEs, which contain “drift” and “diffusion” terms, stochasticity arises from the randomness in the diffusion term, typically modeled as white or colored Gaussian noise. We hypothesize that this diffusion term leads NSDEs to act as a differentially-private (DP) mechanism [33], preserving the privacy of its inputs along its evolution. Here we first formally show that NSDEs are DP-learners. Because of the diffusion term, those models learn “differentially-private” systems of ODEs, achieving privacy without the need of additional mechanisms, such as DP-SGD [23], which alters the training process. Next in §5, we empirically assess the membership risks of NSDEs in comparison to non-stochastic NODE variants and “private” feedforward models trained with DP-SGD.
4.1 Theoretical Analysis
@@
in terms of a drift function ${}f:\Omega\to\mathbb{R}^{n}$ and a diffusion term ${}G:\Omega\to\mathbb{R}^{n\times n}$, where the explicit dependence on parameters $\Theta$ has been dropped for notational convenience. Our privacy bounds will make use of the sensitivity of a function ${}g:\Omega\to\mathbb{R}^{n}$ (c.f. [33]):
@@
Here, we adopt the following natural assumptions from the work by Liu et al. [32], which guarantees the well-posedness of the SDE formulation under consideration.
@@
Theorem 1 (SDEs Are Differentially-Private).
@@
For any privacy budget $10<\varepsilon<$ 1 and failure probability $\delta>0$, the mechanism ${}M(~{}h)=~{}h(T)$, defined by the flow map of the SDE at time $T$, is $(\varepsilon,\delta)$-differentially private, provided that $\sigma\geq\sqrt{2\ln(1.25/\delta)}(TL_{~{}f}/\varepsilon)$ [^2] Note that $\varepsilon$ and $\delta$ here are different from those used in §2.3..
@@
where ${}B_{T}\sim\mathcal{N}\left(~{}0,\sigma^{2}~{}I\right)$ represents additive Gaussian noise and ${}F$ denotes deterministic part of the SDE evolution. Proposition 1 in Appendix shows that this mechanism obeys $(\varepsilon,\delta)$-differential privacy whenever $\varepsilon<1$ and $\sigma\geq\sqrt{2\ln(1.25/\delta)}(S_{~{}F}/\varepsilon)$. The claimed bound on the variance $\sigma$ follows since the sensitivity $S_{~{}F}$ is bounded above due to Lipschitz continuity in ${}f$:
@@
Corollary 1 (NSDE Training is Differentially Private).
Consider target privacy parameters $10\!<\!\varepsilon^{\prime}\!<\!$ 1 and $\delta^{\prime}\!>\!0$ with a maximum iteration count of $K$. Then, NSDE training of the drift ${}f$ in Theorem 1 will be $(\varepsilon^{\prime},K\delta+\delta^{\prime})$-differentially private as long as the variance satisfies $\sigma\!\geq\!4\sqrt{K\ln(1.25/\delta)\ln(1/\delta^{\prime})}(TL/\varepsilon^{\prime})$. Here, $\delta\!>\!0$ is arbitrary and $L=\max_{~{}f}L_{~{}f}$ denotes the maximum Lipschitz constant attainable by ${}f$ during training.
@@
Note that, by the parallel composition principle, mini-batching into disjoint subsets at each iteration does not increase the total privacy leakage, as each individual’s data in the state ${}h$ is accessed only once by the SDE mechanism. Similarly, the composition of the NSDE output with a downstream classifier does not degrade $(\varepsilon,\delta)$-differential privacy guarantees, as these guarantees are immune to post-processing by construction. Consequently, the total privacy leakage in NSDE training is directly proportional to the number of training iterations $K$, corresponding to the number of accesses to the private input ${}h$. The result follows from the Strong Composition Theorem in Dwork et al. [44] [Theorem 3.20, Corollary 3.21], in combination with the variance bound established in Theorem 1. ∎
@@
Note that determining the Lipschitz constant $L$ in Corollary 1 is most easily accomplished by restricting the NSDE architecture to guarantee a particular Lipschitz bound, see, e.g., [45]. Moreover, it should be noted that the privacy guarantees presented here are not necessarily tight in every case and may be improved with more sophisticated privacy accounting mechanisms, such as those in [23, 46].
@@
Corollary 1 guarantees the privacy of NSDE training provided that the diffusion term is constructed to add enough stochastic noise along NSDE solution trajectories. This has the remarkable advantage of privatizing learning without requiring direct intervention in the training process: the network dynamics are already private, therefore no further privacy-preserving modifications are required during the backward pass. In contrast, DP-SGD must use inexact gradient information in order to retain privacy, which can potentially hinder the training process.
@@
TABLE III: Contrasting membership inference risks of NSDEs and NODEs. We evaluate both models with six attacks.
@@
TPR @ 0.1% FPR TPR @ 1% FPR AUC Inference acc. Model Method F-M C-10 C-100 T-I F-M C-10 C-100 T-I F-M C-10 C-100 T-I F-M C-10 C-100 T-I NODE Yeom et al. [3] 0.00% 0.00% 0.01% 0.09% 0.06% 0.00% 0.71% 1.12% 0.518 0.558 0.701 0.645 52.55% 56.59% 70.75% 61.23% Shokri et al. [4] 0.09% 0.04% 0.04% 0.10% 0.90% 0.43% 0.37% 1.01% 0.493 0.448 0.398 0.502 50.15% 50.01% 50.00% 50.21% Song and Mittal [38] 0.00% 0.00% 0.15% 0.12% 0.53% 0.00% 1.66% 1.19% 0.507 0.529 0.498 0.500 50.98% 54.22% 52.48% 50.45% Watson et al. [34] 0.00% 0.00% 0.01% 0.12% 0.04% 0.00% 0.71% 1.17% 0.517 0.558 0.701 0.542 52.04% 56.69% 70.75% 53.24% Carlini et al. [6] 0.19% 1.01% 3.30% 0.24% 1.48% 4.22% 12.11% 1.74% 0.529 0.616 0.782 0.543 52.81% 57.65% 72.88% 52.95% Zarifzadeh et al. [35] 0.67% 1.93% 5.41% 0.26% 3.16% 6.57% 17.58% 2.04% 0.523 0.592 0.797 0.570 53.05% 58.17% 73.37% 55.04% SDENet ($\sigma\!=\!2$) Yeom et al. [3] 0.00% 0.02% 0.05% 0.11% 0.06% 0.92% 1.09% 1.13% 0.526 0.525 0.588 0.557 54.17% 52.44% 56.96% 54.72% Shokri et al. [4] 0.09% 0.09% 0.08% 0.11% 0.94% 0.86% 0.85% 1.03% 0.498 0.484 0.473 0.502 50.14% 50.04% 50.00% 50.40% Song and Mittal [38] 0.00% 0.00% 0.12% 0.12% 0.18% 0.20% 1.34% 1.05% 0.497 0.509 0.498 0.499 50.12% 50.95% 50.84% 50.12% Watson et al. [34] 0.01% 0.02% 0.05% 0.11% 0.18% 0.91% 1.09% 1.12% 0.515 0.525 0.588 0.519 51.02% 52.44% 52.44% 51.91% Carlini et al. [6] 0.10% 0.14% 0.43% 0.12% 0.97% 1.24% 3.00% 1.14% 0.500 0.625 0.612 0.509 50.01% 51.89% 57.64% 50.77% Zarifzadeh et al. [35] 0.27% 0.38% 0.83% 0.16% 1.82% 2.39% 4.60% 1.45% 0.527 0.554 0.646 0.535 51.12% 54.11% 59.25% 52.67%
5 Empirical Evaluation